I am stuck on a quadratic graph question.
y = (x-2)squared,
sorry i can`t do a little 2 so i put squared, i can do other quadratic graphs and the working, but can`t find how to solve this 1

Answer by Pieter(35)

y = (x - 2)²
We have a formula: (a + b)² = a² + 2ab + b² (VERRY IMPORTANT!!!)
or (a - b)² = a² - 2ab + b²
So we get that (x - 2)² = x² - 4x + 4 and you will be able to solve this one.
I find the solution 2 (only one solution because b² - 4ac from the abcD equation is 0: only one solution)
An other way you can get the solutions: find the root of y = x - 2. This wil be (two times due to the exponent) the root of y = (x - 2)² that is because when you search the roots, y = 0 so (x - 2)² = 0 and here the exponent can be left away (sqrt(o) = o anyway)
I hope you understand it, otherwise contact me.
Greets